# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

### v1

@@ 2. Background
Let $L$ be a finitary888“Finitary” means that all principal ideals are finite. For lattices, it is equivalent to $\hat{0}$ exists and the lattice is locally finite. distributive lattice with poset of join-irreducible elements $J_{L}=J(L)$. Then $L$ is the cartesian product of two non-trivial lattices $L=L_{1}\times L_{2}$ iff $J_{L}$ is the disjoint union of two non-empty posets $J_{L}=J_{1}\sqcup J_{2}$, in which case $L_{1}={\operatorname{Ideal}_{\textup{f}}}(J_{1})$ and $L_{2}={\operatorname{Ideal}_{\textup{f}}}(J_{2})$.
@@
Let $J$ be a poset (typically, the poset $J(L)$ of join-irreducibles of a distributive lattice $L$) and $J=J_{1}\sqcup J_{2}$. Then
@@
    If $J$ is cover-modular, then $J_{1}$ and $J_{2}$ are cover-modular. (def. 2.1.)
@@
    If $J$ is unique-cover-modular, then $J_{1}$ and $J_{2}$ are unique-cover-modular. (def. $2.$)
@@
Suppose $L$ is a distributive lattice, $L$ is finitary, and $L$ has finite upward covers. It is not necessary that $J$ has finite upward covers, as is shown by the counterexample shown in fig. 1.
@@
(b) Poset $J=J(L)$ does not have finite upward covers
Figure 1. Counterexample in rem. 2.12
Define $J$ to be the poset with elements $x,y_{0},y_{1},y_{2},\ldots,w_{0},w_{1},w_{2},\ldots$ and the coverings:
@@
  • $x\lessdot y_{i}$ for all $i$, and
  • $w_{i}\lessdot y_{i}$ for all $i$.
Define $L$ to be the lattice $L={\operatorname{Ideal}_{\textup{f}}}(J)$ of all finite ideals of $J$. By Birkhoff’s Representation Theorem[Stan2012a]*Prop. 3.4.3[Birk1967a]*§III.3 Th. 3 $L$ is distributive, $L$ is finitary, and $J$ is poset-isomorphic to the set of join-irreducible elements of $L$. The elements of $L$ are
@@
  • $w_{i},y_{n_{1}},y_{n_{2}},y_{n_{3}},\ldots y_{n_{j}}]=\{x,w_{0},w_{1},w_{2},\ldots,w_{i},y_{n_{1}},y_{n_{2}},y_{n_{3}},\ldots y_{n_{j}}\}$ for all $j\geq 1$, all $n_{1}<n_{2}<n_{3}<\cdots<n_{j}$, and $n_{j}\leq i$ (some non-empty finite set of $y_{\bullet}$ added to a compatible ideal of the preceding type),
Where $a,b,c,\ldots]$” is the ideal of $J$ generated by the elements $a,b,c,\ldots$.
We see that each of the above types of elements of $L$ (ideals of $J$) has only a finite number of ideals that cover it in $L$, that is, contain one more element than it does. So $L$ has finite upward covers. We see that $x$ has all of the $y_{\bullet}$ as upward covers in $J$. So $J$ does not have finite upward covers.
@@ 3. Fomin lattice essentials
The celebrated Robinson–Schensted algorithm[Rob1938a][Schen1961a][Fom1994a][Fom1995a] and its variants that have attracted substantial attention can be constructed[Wor2023c]*sec. 2 from the following machinery. [^9] Though since a Robinson–Schensted variant can be constructed from any dual graded graph, it is not clear why the rest of the broader class has not been as productive. See sec. 8(9). The core of the machinery is a “Fomin lattice” which satisfies the properties discussed in para. 2 and tersely defined below.
@@
      a weighting or weight $w(\bullet\lessdot\bullet)$, which is a function from coverings (or prime quotients), pairs $x\lessdot y$ of elements of $L$, to $V$, for which $L$ is a weighted-differential lattice[Stan1988a][Fom1994a]:
@@
      for which $w$ is constant on projectivity classes[Birk1967a]*Ch. V sec. 5 of coverings or equivalently, on coverings which are cover-projective: if $x\wedge y\lessdot x$ and $y\lessdot x\vee y$, then $w(x\wedge y\lessdot x)=w(y\lessdot x\vee y)$ and $w(x\wedge y\lessdot y)=w(x\lessdot x\vee y)$.
@@
    A McLarnan map [^11] We name the correspondence for McLarnan, who seems to have been the first to emphasize[McLar1986a]*p. 22 the arbitrariness of the choice of the correspondence and thus that there is no “natural” correspondence. The set of possible correspondences for a particular Fomin lattice is usually uncountable. We choose “map” because of its consonance with “McLarnan”. or $R$-correspondence $M$, which is a realization of the weighted-differential condition:
@@
    Depending on the degree of abstraction of the discussion, the requirements for the set of values $V$ varies. The minimum requirement is that $V$ be a module over a unitial ring of scalars $S$. Usually we will be unconcerned with the details of $S$ other than it necessarily contains a homomorphic image of $\mathbb{Z}$ and thus we can multiply elements of $V$ by integer coefficients. If $V=\mathbb{Z}$ (which implies $\mathbb{Z}$ is isomorphically embedded in $S$), $r>0$, and all of the values of $w$ are positive, then the weighting is defined to be admissible, which is required to construct a Robinson–Schensted algorithm.
@@
Proof.
In light of the above lemma, we will abuse language by using “$w$ is admissible” or “the lattice is admissible” to mean that the weighting is admissible, that is, $V$ is $\mathbb{Z}$, $r>0$, and all of the values of $w$ are positive.
@@
If a lattice $L$ forms a Fomin lattice with two pairs of differential degrees and weightings $(r,w)$ and $(r^{\prime},w^{\prime})$, then $L$ forms a Fomin lattice with any integer linear combination of the two: the differential degree $r^{\prime\prime}=\alpha r+\beta r^{\prime}$ and weighting $w^{\prime\prime}(x\lessdot y)=\alpha w(x\lessdot y)+\beta w^{\prime}(x\lessdot y)$ for any $\alpha,\beta\in\mathbb{Z}$. Thus, the set of pairs of degrees and weightings for $L$ forms a subspace of an infinite-dimensional module over $S$, the subspace of solutions of the set of linear equations (7).
If $(r,w)$ and $(r^{\prime},w^{\prime})$ are admissible, $\alpha\geq 0$, $\beta\geq 0$, and not both $\alpha=\beta=0$, then $(r^{\prime\prime},w^{\prime\prime})$ is admissible.
Lemma 3.4.
Because $L$ is modular and $w$ satisfies the projectivity condition (A)(4), the Kurosh–Ore theorem[Dil1961a]*Th. 3.3 shows that $L$ is graded and that the multiset of the weights of the coverings composing a path from some $x$ to some $y\geq x$ in $L$ depends only on $x$ and $y$ and not on the particular path.
@@
Given this machinery, the general Robinson–Schensted insertion algorithm can be constructed using Fomin’s growth diagram construction provided the weighting is admissible. The set of permutations of $\{1,\ldots,n\}$ whose elements are $r$-colored can be bijectively mapped to the set
$\{(P,Q,c)\,|\,$ there exists $x\in L$ of rank $n$ for which
@@
where the particular bijection depends on the McLarnan map that is chosen. (We use “path” rather than the more usual “saturated chain” to emphasize the importance of the coverings composed of adjacent elements of the chain, which bear the weights.)
@@
Each element $x$ of $L$ with rank $n$ is called a diagram of size $n$. Each path from $\hat{0}$ to $x$ is called a tableau of shape $x$ and thus size $n$. We define $f(x)$ as the number of paths from $\hat{0}$ to $x$, that is, the number of tableaux of shape $x$.
@@ 4. Auxiliary result
Let $L$ be a locally finite modular lattice and let $S$ be a set of coverings in $L$ that is closed under cover-projectivity, that is, if $x\lessdot y$ is in $S$ and it is cover-projective to $x^{\prime}\lessdot y^{\prime}$, then $x^{\prime}\lessdot y^{\prime}$ is in $S$. Since $L$ is locally finite and modular, this is equivalent to $S$ is the union of a set of projectivity classes of coverings in $L$.
@@
Since $L$ is modular and locally finite, the interval $[x,y]$ is modular and finite. Choose any projectivity class $C$ of coverings in $L$. The Kurosh–Ore theorem can be applied $[x,y]$ to show that one path contains a covering in $C$ iff the other path contains a covering in $C$. The lemma then follows. ∎
@@
We define $x\sim y$ iff no covering in a path from $x\wedge y$ to $x\vee y$ is in $S$.
@@
By lem. 4.3, no covering in any path from $x$ to $z$ is in $S$. We can choose such a path from $x$ to $y$ and then from $y$ to $z$. Since no covering in that path is in $S$, no covering from $x$ to $y$ or from $y$ to $z$ is in $S$. Then by lem. 4.3, $x\sim y$ and $y\sim z$. ∎
@@
This follows directly from the diamond isomorphism theorem and lem. 4.3. ∎
@@
If $\in L, xyx\sim y$ implies $\wedge y\sim x\vee y, xx$$\sim x\vee y\sim y$, and $x\sim x\wedge y\sim y$.
@@
Given $x\sim x\vee y$, by lem. 4.6, $x\wedge y\sim y$. Since $x\wedge y\leq y\leq x\vee y$, $x\wedge y\sim y$, and $y\sim x\vee y$, By lem. 4.5, $x\wedge y\sim x\vee y$ and so $x\sim y$. ∎
@@
We know $x\sim y$ and $y\sim z$. By lem. 4.7, $x\sim x\vee y$, $x\vee y\sim y$, $y\sim y\vee z$, and $y\vee z\sim z$. Define $w=(x\vee y)\wedge(y\vee z)$. Of necessity, $y\leq w\leq x\vee y$ and $y\leq w\leq y\vee z$. By lem. 4.4, $w\sim x\vee y$ and $w\sim y\vee z$. By lem. 4.8, $x\vee y\sim y\vee z$ and by lem. 4.7, $x\vee y\sim(x\vee y)\vee(y\vee z)=x\vee y\vee z$. Since $x\leq x\vee y\leq x\vee y\vee z$, by lem. 4.5, $x\sim x\vee y\vee z$. Since $x\leq x\vee z\leq x\vee y\vee z$, by lem. 4.4, $x\sim x\vee z$. Similarly, $z\sim x\vee z$. Then by lem.4.8, $x\sim z$. ∎
@@
By lem. 4.9, $y\sim z$, and by lem. 4.7, $y\sim y\wedge z$. Then by lem. 4.5, $x\sim y\wedge z$. Dually, $x\sim y\vee z$. ∎
@@
Given $x,y,z,w\in L$ with $y\leq z\leq w$, if $x\sim y$ and $x\sim w$, then $x\sim z$.
@@
We define $L_{/Sz}$ to be $\{x\in L\,|\,x\sim z\}$ (where $\sim$ is determined by $S$).
@@
We apply this construction to a Fomin lattice that has a non-admissible weighting.
@@
Given a Fomin lattice $L$ with weighting $w$ and $\hat{0}$, we define its truncation $L_{/}$ to be $L_{/S\hat{0}}$, where $S=\{x\lessdot y\,|\,w(x\lessdot y)=0\}$, the set of coverings with weight 0. We define $w_{/}(x\lessdot y)$ for any $x,y\in L/$ to be equal to $w(x\lessdot y)$ (for $L$).
@@
Given a Fomin lattice $L$ with $\hat{0}$, its truncation $L_{/}$ is a Fomin lattice.
@@
All of the requirements for $L_{/}$ to be a Fomin lattice (def. 3.1) are trivially inherited from $L$ except for (7). But by construction, (7) for a particular $x\in L_{/}$ contains a subset of the terms of (7) for $x\in L$, and the terms of the latter that are not in the former are ones for which $w$ has the value 0. Thus (7) is true for all $x\in L_{/}$. ∎
@@
Given a distributive Fomin lattice $L$ with $\hat{0}$ with points $J(L)$, its truncation $L_{/}$ is isomorphic to ${\operatorname{Ideal}_{\textup{f}}}(J_{/})$ where $J_{/}$ is the subposet $\{j\in J\,|\,\textup{not }(\exists k)(k\leq j\textup{ and }w(j)=0)\}$.
@@
Given a Fomin lattice $L$ with $\hat{0}$, if for every covering $x\lessdot y$ for which $w(x\lessdot y)<0$, there exists a $z,w$ for which $z\leq w\leq x$ and $w(z\leq w)=0$, then $L_{/}$ is a Fomin lattice with an admissible weighting.
@@
Given a distributive Fomin lattice $L$ with $\hat{0}$ with points $J(L)$, if for every point $j$ for which $w(j)<0$ then there exists a $k$ for which $k\leq j$ and $w(k)=0$, then $L_{/}$ is a Fomin lattice with an admissible weighting.
Of course, truncation could be done using a different element $z\in L$ than $\hat{0}$, but if $L$ has hatzerohatzero, then any sublattice of $L$ has a minimum element.
The only known application of truncation is when $L$ is $\mathbb{Y}$, $J$ is the quadrant, and the weighting on $J$ is $w((i,j)=\alpha+\beta i+\gamma j$.[Fom1994a]*Exam. 2.2.7 If $\alpha=r$, $\beta=r/k$, and $\gamma=-r/k$ for some $k$ (positive or negative) that divides $r$, then $L_{/}$ is the $k$-strip (if $k$ is positive) or the dual of the $k$-strip (if $k$ is negative).

### v2

@@ 2. Background
Let $L$ be a finitary777“Finitary” means that all principal ideals are finite. For lattices, it is equivalent to $\hat{0}$ exists and the lattice is locally finite. distributive lattice with poset of join-irreducible elements $P=J(L)$. Then $L$ is the cartesian product of two non-trivial lattices $L=L_{1}\times L_{2}$ iff $P$ is the disjoint union of two non-empty posets $P=P_{1}\sqcup P_{2}$, in which case $L_{1}={\operatorname{Ideal}_{\textup{f}}}(P_{1})$ and $L_{2}={\operatorname{Ideal}_{\textup{f}}}(P_{2})$. [^8] We define ${\operatorname{Ideal}_{\textup{f}}}(P)$, where $P$ is a poset, to be the lattice of finite lower order ideals of $P$.
@@
Let $P$ be a poset and $P=P_{1}\sqcup P_{2}$. Then
@@
    If $P$ is cover-modular, then $P_{1}$ and $P_{2}$ are cover-modular. (def. 2.1.)
@@
    If $P$ is unique-cover-modular, then $P_{1}$ and $P_{2}$ are unique-cover-modular. (def. $2.$)
@@
Suppose $L$ is a distributive lattice, $L$ is finitary, and $L$ has finite upward covers. It is not necessary that $P$ has finite upward covers, as is shown by the counterexample shown in fig. 5.
@@
(b) Poset $P=J(L)$ does not have finite upward covers
Figure 5. Counterexample in rem. 2.13
Define $P$ to be the poset with elements $x,y_{0},y_{1},y_{2},\ldots,w_{0},w_{1},w_{2},\ldots$ and the coverings:
@@
  • $x\lessdot y_{i}$ for all $i\geq 0$, and
  • $w_{i}\lessdot y_{i}$ for all $i\geq 0$.
Define $L$ to be the lattice $L={\operatorname{Ideal}_{\textup{f}}}(P)$ of all finite ideals of $P$. By Birkhoff’s Representation Theorem[Birk1967a]*§III.3 Th. 3[Stan2012a]*Prop. 3.4.3 $L$ is distributive, $L$ is finitary, and $P$ is poset-isomorphic to the set of join-irreducible elements of $L$. The elements of $L$ are (where $a,b,c,\ldots]$ is the ideal of $P$ generated by the elements $a,b,c,\ldots$):
@@
  • $w_{i},y_{n_{1}},y_{n_{2}},y_{n_{3}},\ldots y_{n_{j}}]=\{x,w_{0},w_{1},w_{2},\ldots,w_{i},y_{n_{1}},y_{n_{2}},y_{n_{3}},\ldots y_{n_{j}}\}$ for all $i\geq 0$, all $j\geq 1$, and all $0\leq n_{1}<n_{2}<n_{3}<\cdots<n_{j}\leq i$ (some non-empty finite set of $y_{\bullet}$ added to a compatible ideal of the preceding type),
We see that each of the above types of elements of $L$ (ideals of $P$) has only a finite number of ideals that cover it in $L$, that is, contain one more element than it does. So $L$ has finite upward covers. We see that $x$ has all of the $y_{\bullet}$ as upward covers in $P$. So $P$ does not have finite upward covers.
@@ 3. Fomin lattice essentials
The celebrated Robinson–Schensted algorithm[Rob1938a][Schen1961a][Fom1994a][Fom1995a] and its variants that have attracted substantial attention can be constructed[Wor2023c]*sec. 2 from the following machinery. [^9] Though since a Robinson–Schensted variant can be constructed from any DGG, it is not clear why DGGs that are not Fomin lattices have not been as productive. See item (10) in sec. 8. The core of the machinery is a “Fomin lattice” which satisfies properties (A)(G) of para. 2, summarized here:
@@
      a weighting $w(\bullet\lessdot\bullet)$, which is a function from coverings (or prime quotients), pairs $x\lessdot y$ of elements of $L$, to $V$, for which $L$ is a weighted-differential lattice[Stan1990a]*sec. 3[Fom1994a]*sec 2.2:
@@
      for which $w$ is projective-constant, that is, constant on projectivity classes[Birk1967a]*§I.7 Def. of coverings, or equivalently, on coverings which are cover-projective: if $x\vee y$ covers $x$ and $y$ and $x\wedge y$ is covered by $x$ and $y$, then $w(x\wedge y\lessdot x)=w(y\lessdot x\vee y)$ and $w(x\wedge y\lessdot y)=w(x\lessdot x\vee y)$.
@@
    A McLarnan map [^11] We name the correspondence for McLarnan, who seems to have been the first to emphasize[McLar1986a]*Sec. 1 p. 22 the arbitrariness of the choice of the correspondence and thus that there is no “natural” correspondence. The set of possible correspondences for a particular Fomin lattice is usually uncountable. We choose “map” because of its consonance with “McLarnan”. or $R$-correspondence $M$, which is a realization of the weighted-differential condition:
@@
    Depending on the degree of abstraction of the discussion, the requirements for the set of values $V$ varies. The minimum requirement is that $V$ be a module over a unitial ring of scalars $S$. Usually we will be unconcerned with the details of $S$ other than it necessarily contains a homomorphic image of $\mathbb{Z}$ and thus we can multiply elements of $V$ by integer coefficients. If $V=\mathbb{Z}$ (which implies $\mathbb{Z}$ is isomorphically embedded in $S$), $r>0$, and all of the values of $w$ are positive, then the weighting is defined to be positive, which is required to construct a Robinson–Schensted algorithm.
@@
If a lattice $L$ forms a Fomin lattice with two pairs of differential degrees and weightings $(r,w)$ and $(r^{\prime},w^{\prime})$, then $L$ forms a Fomin lattice with any linear combination of the two: the differential degree $r^{\prime\prime}=\alpha r+\beta r^{\prime}$ and weighting $w^{\prime\prime}(x\lessdot y)=\alpha w(x\lessdot y)+\beta w^{\prime}(x\lessdot y)$ for any $\alpha,\beta\in S$. Thus, the set of pairs of degrees and weightings for $L$ forms a subspace of an infinite-dimensional module over $S$, the subspace of solutions of the set of linear equations (7).
If $(r,w)$ and $(r^{\prime},w^{\prime})$ are positive, $\alpha\geq 0$, $\beta\geq 0$, and not both $\alpha$ and $\beta$ are 0, then $(r^{\prime\prime},w^{\prime\prime})$ is positive.
@@
Because $L$ is modular and $w$ satisfies the projectivity condition $(A)$, the relevant version of the Jordan–Hölder theorem[Birk1967a]*§III.7 Th. 9 Cor.[Nat2017a]*Th. 9.7 shows that $L$ is graded and that the multiset of the weights of the coverings composing a path from some $x$ to some $y\geq x$ in $L$ depends only on $x$ and $y$ and not on the particular path.
@@
Given this machinery, the general Robinson–Schensted insertion algorithm can be constructed using Fomin’s growth diagram construction provided the weighting is positive. The set of permutations of $\{1,\ldots,n\}$ whose elements are $r$-colored can be bijectively mapped to the set
$\{(P,Q,c)\,|\,$ there exists $x\in L$ of grade $n$ for which
@@
where the particular bijection depends on the McLarnan map that is chosen. [^12] We use “path” rather than the more usual “saturated chain” to emphasize the importance of the coverings composed of adjacent elements of the chain, which bear the weights.
@@
Each element $x$ of $L$ with grade $n$ is called a diagram of size $n$. Each path from $\hat{0}$ to $x$ is called a tableau of shape $x$ and thus size $n$. We define $f(x)$ as the number of paths from $\hat{0}$ to $x$, that is, the number of tableaux of shape $x$.
@@ 4. Auxiliary result
Let $L$ be a locally finite modular lattice and let $S$ be a set of coverings in $L$ that is closed under cover-projectivity, that is, if $x\lessdot y$ is in $S$ and it is cover-projective to $x^{\prime}\lessdot y^{\prime}$, then $x^{\prime}\lessdot y^{\prime}$ is in $S$. Since $L$ is locally finite and modular, this is equivalent to that $S$ is the union of a set of projectivity classes of coverings in $L$.
@@
Since $L$ is modular and locally finite, the interval $[x,y]$ is modular and finite. Choose any projectivity class $C$ of coverings in $L$. The relevant version of the Jordan–Hölder theorem[Birk1967a]*§III.7 Th. 9 Cor.[Nat2017a]*Th. 9.7 can be applied $[x,y]$ to show that one path contains a covering in $C$ iff the other path contains a covering in $C$. The lemma then follows. ∎
@@
We define $x\sim_{S}y$ (or $x\sim y$ if, as usual, $S$ is understood) iff no covering in a path from $x\wedge y$ to $x\vee y$ is in $S$.
@@
By lem. 4.3, no covering in any path from $x$ to $z$ is in $S$. We can choose such a path from $x$ to $y$ and from there to $z$. Since no covering in that path is in $S$, no covering from $x$ to $y$ or from $y$ to $z$ is in $S$. Then by lem. 4.3, $x\sim y$ and $y\sim z$. ∎
@@
This follows from the diamond isomorphism theorem and lem. 4.3. ∎
@@
If $\in L, xyx\sim y$ implies $x\wedge y\sim x\vee y$, $x\wedge y\sim x\sim x\vee y$, and $x\wedge y\sim y\sim x\vee y$.
@@
Given $x\sim x\vee y$, by lem. 4.6, $x\wedge y\sim y$. Since $x\wedge y\leq y\leq x\vee y$, $x\wedge y\sim y$, and $y\sim x\vee y$, By lem. 4.5, $x\wedge y\sim x\vee y$ and so $x\sim y$. The dual is proved dually. ∎
@@
The construction for this proof is shown in fig. 6.
Figure 6. Construction for the proof of lem. 4.9.
We know $x\sim y$ and $y\sim z$. By lem. 4.7, $x\sim x\vee y$, $x\vee y\sim y$, $y\sim y\vee z$, and $y\vee z\sim z$. Define $v=(x\vee y)\wedge(y\vee z)$. Of necessity, $y\leq v\leq x\vee y$ and $y\leq v\leq y\vee z$. By lem. 4.4, $v\sim x\vee y$ and $v\sim y\vee z$. By lem. 4.8, $x\vee y\sim y\vee z$ and by lem. 4.7, $x\vee y\sim(x\vee y)\vee(y\vee z)=x\vee y\vee z$. Since $x\leq x\vee y\leq x\vee y\vee z$, by lem. 4.5, $x\sim x\vee y\vee z$. Since $x\leq x\vee z\leq x\vee y\vee z$, by lem. 4.4, $x\sim x\vee z$. Similarly, $z\sim x\vee z$. Then by lem. 4.8, $x\sim z$. ∎
@@
By lem. 4.9, $y\sim z$, and by lem. 4.7, $y\sim y\wedge z$. Then by lem. 4.9, $x\sim y\wedge z$. Dually, $x\sim y\vee z$. ∎
@@
Given $x,y,z,v\in L$ with $y\leq z\leq v$, if $x\sim y$ and $x\sim v$, then $x\sim z$.
@@
We define $L_{/Sz}$ to be the equivalence class of $\sim_{S}$ containing $z$, which is $\{x\in L\,|\,x\sim_{S}z\}$.
@@
We now apply this construction to a Fomin lattice that has a non-positive weighting.
@@
Given a Fomin lattice $L$ with weighting $w$ and $\hat{0}$, we define its truncation $L_{/}$ to be $L_{/S\hat{0}}$, where $S=\{x\lessdot y\,|\,w(x\lessdot y)=0\}$, the set of coverings with weight 0. We define the weighting $w_{/}$ on $L_{/}$ to be $w$ restricted to $L_{/}$.
@@
Given a Fomin lattice $L$ with $\hat{0}$, its truncation $L_{/}$ is a Fomin lattice with (the same) $\hat{0}$.
@@
All of the requirements for $L_{/}$ to be a Fomin lattice (def. 3.1) are trivially inherited from $L$ except for (A). But by construction, equation (7) for a particular $x\in L_{/}$ contains a subset of the terms of (7) for $x\in L$, and the terms of the latter that are not in the former are ones for which $w$ has the value 0. Thus (7) is true for all $x\in L_{/}$. ∎
@@
Given a finitary distributive Fomin lattice $L$ with $\hat{0}$ with points $J(L)$, its truncation $L_{/}$ is isomorphic to ${\operatorname{Ideal}_{\textup{f}}}(P_{/})$ where $P_{/}$ is the subposet $\{j\in J(L)\,|\,\textup{not }(\exists k)(k\leq j\textup{ and }w(j)=0)\}$.
@@
Given a Fomin lattice $L$ with $\hat{0}$, if its set of values is $\mathbb{Z}$, $r>0$, and for every covering $x\lessdot y$ for which $w(x\lessdot y)<0$, there exists a $z,v$ for which $z\lessdot v\leq x$ and $w(z\lessdot v)=0$, then $L_{/}$ is a Fomin lattice with $\hat{0}$ and a positive weighting.
Proof.
By th. 4.16, $L_{/}$ is a Fomin lattice with $\hat{0}$. It remains to prove that its weighting $w_{/}$ is has only positive values on $L_{/}$. Assume to the contrary that $x\lessdot y$ in $L_{/}$ and $w_{/}(x\lessdot y)\leq 0$. If $w_{/}(x\lessdot y)=0$, then by the construction of $L_{/}$ and lem. 4.3, $y\not\in L_{/}$. If on the other hand $w_{/}(x\lessdot y)<0$, then by hypothesis there is $\hat{0}\leq z\lessdot v\leq y$ with $w_{/}(z\lessdot v)=0$, and similarly $v\not\in L_{/}$, implying $y\not\in Z_{/}$. ∎
@@
Given a distributive Fomin lattice $L$ with $\hat{0}$ with points $P(L)$, if for every point $j$ for which $w(j)<0$ then there exists a point $k$ for which $k\leq j$ and $w(k)=0$, then $L_{/}$ is a Fomin lattice with a positive weighting.
Of course, this construction of a smaller lattice that is an positive Fomin lattice could be done by choosing the component containing a different element $z\in L$ than $\hat{0}$, since if $L$ has zerohatzerohat, then any sublattice of $L$ has a minimum element.
The only known application of truncation is when $L$ is $\mathbb{Y}$, $P=J(L)$ is the quadrant, and the weighting on $P$ is $w((i,j)=\alpha+\beta i+\gamma j$.[Fom1994a]*Exam. 2.2.7 If $\alpha=r$, $\beta=r/k$, and $\gamma=-r/k$ for some nonzero integer $k$ that divides $r$, then $L_{/}$ is $\mathbb{Y}_{k}$ with the weights multiplied by $r/k$ (if $k$ is positive) or its left-right reversal (if $k$ is negative).